# ChOiRe: Characterizing and Predicting Human Opinions with Chain of Opinion Reasoning

### v3

@@ Abstract
Warning: This paper includes examples that may be deemed sensitive or offensive.
Aligning language models (LMs) with human opinion is challenging yet vital to enhance their grasp of human values, preferences, and beliefs. We present ChOiRe [^1] Codes and data will be available at https://github.com/dxlong2000/ChOiRe., a four-step framework to predict human opinion which differentially models the user’s explicit personae (i.e. demographic or ideological attributes) that are manually declared, and implicit personae inferred from user historical opinions. ChOiRe consists of (i) an LM analyzing the user’s explicit personae to filter out irrelevant attributes; (ii) the LM ranking the implicit persona opinions into a preferential list; (iii) Chain-of-Opinion (CoO) reasoning, where the LM sequentially analyzes the explicit personae and the most relevant implicit personae to perform opinion prediction; (iv) and where ChOiRe executes Step (iii)’s CoO multiple times with increasingly larger lists of implicit personae to overcome insufficient personae information to infer a final result. ChOiRe achieves new state-of-the-art effectiveness with limited inference calls, improving previous techniques significantly by 3.22%. We also show that ChOiRe’s Steps (i) and (ii) can significantly better fine-tune opinion-aligned models, by up to 18.44%.
@@ 1 Introduction
Language models (LMs) are becoming indispensable tools, serving in various roles such as dialogue agents OpenAI (2022); Google (2022), data analysts Wang et al. (2023a); Cheng et al. (2023), and decision support Ye et al. (2023). LMs also demonstrate the capability to model distinct opinions which influence response generation on input queries Bai et al. (2022); Glaese et al. (2022); Santurkar et al. (2023). Unfortunately, the opinions modeled by language models are shaped by the extensive training and feedback data, which are themselves influenced by countless human perspectives, making them inherently challenging to model. As human–AI interactions become common, it becomes imperative to align models with human opinion to meet individual and personalized expectations.
Alignment frameworks such as RLHF Christiano et al. (2017); Ouyang et al. (2022) have been devised to train and fine-tune LMs. Nonetheless, applying them to align large language models (LLMs) with human opinions is difficult because they require both significant computational resources and high-quality supervised feedback data, which is itself challenging to collect. Therefore, prompt-based opinion alignment through persona has been studied as an easy-to-deploy and easy-to-combine resource-parsimonious alternative Perez et al. (2023); Simmons (2023); Santurkar et al. (2023); Deshpande et al. (2023).
However, even when aligning with well-represented groups, persona-based methods result in low steerability Santurkar et al. (2023). This raises significant concerns when using LMs to model for individual users. Since individuals hold nuanced opinions that evolve and are influenced by situational factors, it is more challenging to achieve individual opinion LM alignment. Hwang et al. (2023) find significant opinion variations among individuals sharing the same demographics, exposing flaws in current group-focused alignment. They argue for individualized models, introducing an approach (detailed below) integrating user’s demographic and ideological attributes, (we term as explicit personae) and historical opinions (implicit personae) into the prompt for opinion prediction.
While Hwang et al. (2023)’s naïve integration achieves promising results, we argue that it suffers from a few key limitations. First, it employs all explicit personae. However, we contend that only a subset is needed for accurate opinion prediction; including non-relevant personae may act as noise, harming predictive performance. Second, Hwang et al. (2023) utilize the top-$K$ semantically similar opinions with the question (here termed top-$K$ implicit personae). This approach is inefficient, as such similar opinions may not offer the most valuable information for opinion prediction. Additionally, our empirical experiments suggest that LMs may lack sufficient personae evidence with a fixed $K$ — dynamically adjusting $K$ per task can overcome such deficiencies. Finally, while Chain-of-Thought (CoT; Wei et al. 2022; Kojima et al. 2022) enables LMs to perform multi-step reasoning tasks effectively, we find that the naïve application of CoT does not help modern LLMs like ChatGPT with opinion alignment [^2] We validate all our claims on these limitations in Sections E.1, 6.1, 4 and 5, respectively..
To address the above challenges, we propose ChOiRe [^3] Chain of Opinion Reasoning, pronounced as the English word “choir”.(fig. 1), a four-step solution for opinion prediction leveraging LLMs’ strong data evaluation and analytic capabilities Wang et al. (2023a); Cheng et al. (2023). First, an LLM analyzes a target user’s explicit personae to discard irrelevant ones. Second, the LLM ranks implicit persona opinions in order of usefulness, selecting the top-$K$ as the most valuable. This surpasses the constraint of using semantic similarity scores. Third, we introduce Chain-of-Opinion (CoO), a designed variant of CoT that allows the LLM to explain and analyze selected explicit personae and top-$K$ implicit personae sequentially. ChOiRe applies self-consistency over CoO to provision the appropriate amount of user information for opinion inference.
ChOiRe achieves new state-of-the-art (SOTA) in opinion alignment effectiveness and reliability, while using a limited inference budget. We conduct a thorough analysis to verify our hypotheses concerning explicit and implicit personae and defend our Chain-of-Opinion reasoning methodology. Our analysis further reveals that ChOiRe’s first two steps in handling explicit and implicit personae also helps to fine-tune opinion-aligned models.
@@ 2 Related Work
Aligning LMs with Humans.
Aligning language models with human behaviour is a recent area of study as alignment can increase user experience satisfaction and utility Wang et al. (2023c). One line of work develops prompting techniques with user demographic information (e.g., political identity) to encourage LMs to output human-like responses. Argyle et al. (2023) show that by properly conditioning LMs with targeted identity and personality profiles, it is possible to produce biased outputs that strongly correlate with human responses. Furthermore, Simmons (2023) claims that LLMs are moral mimics: by giving models a political identity, they produce texts mirroring the associated moral biases. Despite recent advances, Santurkar et al. (2023) discovered that LMs align poorly with human opinions, as evidenced by model performance on public opinion polls. Hwang et al. (2023) recently propose to incorporate explicit and implicit personae to predict human opinions in new contexts. In section 1, we argue that this naïve strategy is suboptimal. ChOiRe overcomes these limitations.

### v5

# Aligning Large Language Models with Human Opinions through Persona Selection and Value–Belief–Norm Reasoning

@@ Abstract
Reasoning and predicting human opinions with large language models (LLMs) is essential yet challenging. Current methods employ role-playing with personae but face two major issues: LLMs are sensitive to even a single irrelevant persona, changing up to 30% of the predictions; and LLMs fail to reason strategically over personae. We propose Chain-of-Opinion (COO [^1] Our codes and data will be made publicly available here.), a simple four-step solution modeling which and how to reason with personae, inspired by the Value–Belief–Norm (VBN) theory. COO differentiates between explicit personae (demographics and ideology) and implicit personae (historical opinions), involves: (1) filtering irrelevant attributes from explicit personae; (2) ranking implicit personae into a preferential list for selecting top-k; (3) applying novel VBN reasoning to extract user environmental and personal value, belief, and norm variables for accurate and reliable predictions; and (4) iterating VBN reasoning with progressively larger lists of implicit personae to handle potential persona insufficiency. COO efficiently achieves new state-of-the-art opinion prediction via prompting with only 5 inference calls, improving prior techniques by up to 4%. Notably, fine-tuning LMs with COO’s data results in significantly better opinion-aligned models, by up to 23%.
Aligning Large Language Models with Human Opinions through
Persona Selection and Value–Belief–Norm Reasoning
Do Xuan Long1,2, Kenji Kawaguchi1, Min-Yen Kan1, Nancy F. Chen2 1National University of Singapore, 2Institute for Infocomm Research (I2R), A*STAR xuanlong.do@u.nus.edu, {kenji,knmnyn}@nus.edu.sg, nfychen@i2r.a-star.edu.sg
@@ 1 Introduction
Pre-trained large language models (LLMs) are becoming indispensable tools, serving in various assistant roles such as dialogue agents OpenAI (2022); Google (2022) and data analysts Cheng et al. (2023). Notably, they demonstrate the capability to model distinct opinions that influence response generation Bai et al. (2022); Glaese et al. (2022); Santurkar et al. (2023). Unfortunately, their opinions are shaped by extensive training data, which are themselves influenced by countless human perspectives and thus challenging to comprehend. As human–AI interactions grow, instructing models to reason in alignment with human opinions is crucial for effective personalization.
Although fine-tuning alignment methods such as RLHF Christiano et al. (2017); Ouyang et al. (2022) are widely employed, their application to personalized opinions remains challenging due to the significant compute and data required. Prompt-based role-playing frameworks using personae have emerged as alternatives. Early work in this direction focused on aligning models with social groups rather than individuals Perez et al. (2023); Santurkar et al. (2023). However, Santurkar et al. (2023) found that simple persona-based prompting exhibits low steerability, even for well-represented groups. Further, Hwang et al. (2023) revealed significant opinion variation among individuals with similar demographics, underscoring the challenge of aligning LLMs to individual opinions. They subsequently proposed a naïve solution to model individuals by incorporating the user’s demographics and ideology (we term these explicit personae), alongside their historical opinions (implicit personae).
While naïvely incorporating explicit and implicit personae into the prompt shows promise for individualization Santurkar et al. (2023); Hwang et al. (2023), this approach is suboptimal. Both persona types include ones that may be irrelevant to the opinion of interest, such as the ‘‘Citizenship’’ for the ‘‘Gun’’ question in Fig. 1. This poses a problem as we observe that LLMs are highly sensitive to such noise (detailed in Section 3), highlighting the task of relevant persona selection as important yet unsolved. Moreover, effective reasoning over explicit and implicit personae for opinion prediction is challenging: we find that Chain-of-Thought Wei et al. (2022); Kojima et al. (2022) unexpectedly fails to improve opinion prediction in LLMs, due to reasoning inconsistencies.
In line with the Value-Belief-Norm (VBN) theory Stern et al. (1999) which asserts that values, beliefs, and norms influence human behavior distinctly, we argue that the explicit and implicit personae should be processed and utilized differently. The explicit offers clear insights into user environmental values (binary relevancy), while the implicit reveals nuanced and context-specific beliefs and norms shaped by personal experiences (complex relevancy). Building on this, we introduce Chain-of-Opinion (COO, Fig. 1), a novel four-step framework that optionally leverages both persona types to address the above challenges: (1) an LLM analyzes explicit personae to discard irrelevant ones; (2) the LLM ranks implicit persona opinions by usefulness and selects the top-$K$; (3) VBN reasoning where the LLM generates high-level environmental values from selected explicit personae and nuanced individual beliefs and norms from top-K implicit personae before deriving a prediction; and (4) COO applies VBN reasoning using varying $K$ of implicit personae from step (2) to prevent the model from refusing to answer due to insufficient personae.
COO achieves state-of-the-art opinion prediction with persona-based prompting by just five inference calls (Section C.1). Moreover, fine-tuning with data from COO’s steps (1–3) enhances LMs by up to 23%, resulting in Flan-T5 base model Chung et al. (2024) comparable with GPT-4 OpenAI (2023b). COO is highly generalizable in scenarios with missing personae (Section 6.1), and its four steps can motivate and be applicable to other personalized tasks involving explicit personae and user historical views (Section B.8).
@@ 2 Related Work
LLM role-play with personae.
Aligning language models with human behavior via personae is a growing study area. Such alignment increases user satisfaction and personalization Wang et al. (2023c); Chen et al. (2024). One line of work develops prompting techniques with user demographics, encouraging LLMs to output human-like responses. Argyle et al. (2023) showed that by properly conditioning LLMs with targeted identity profiles, they produce biased outputs that strongly correlate with human responses. Furthermore, Simmons (2023) claimed that LLMs are moral mimics: by giving models a political identity, they produce texts mirroring the associated moral biases. Nevertheless, Santurkar et al. (2023); Hwang et al. (2023) discovered that LLMs align poorly with human opinions, as evidenced by model performance using explicit and implicit personae on public opinion polls. We argue that this strategy is suboptimal (Section 1) due to noisy personae and the inefficient reasoning strategy employed. COO overcomes these limitations.